# Rethinking Disentanglement under Dependent Factors of Variation

### v2

@@ Appendix A Proofs of Theorems
Let $\bm{y}=\{y_{l}\}_{l=1}^{n}$ be some factors and $\bm{z}=\{z_{j}\}_{j=1}^{m}$ a representation. If $z_{j}$ is a minimal representation of $y_{i}$, then $z_{j}$ is factors-invariant for $y_{i}$ and nuisances-invariant. Equivalently, we have that:
$(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longleftrightarrow(\bm{\tilde{y}}_{i}\leftrightarrow y_{i}\leftrightarrow z_{j})\,\land\,(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})$
@@
Second, we demonstrate that $(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longrightarrow(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})$: Given this Markov chain, we know from the DPI that $I(z_{j};\bm{x})\leq I(z_{j};y_{i})$. On the other hand, since $z_{j}$ is a representation of $\bm{x}$, we have the Markov chain $(\bm{n},y_{i})\leftrightarrow\bm{x}\leftrightarrow z_{j}$. Equivalently, we have that $I(z_{j};\bm{n},y_{i})\leq I(z_{j};\bm{x})$. The chain rule of mutual information tells us that $I(z_{j};\bm{n}|y_{i})=I(z_{j};\bm{n},y_{i})-I(z_{j}|y_{i})$. According to the above two points, we are left with $I(z_{j};\bm{n}|y_{i})\leq I(z_{j};\bm{x})-I(z_{j}|y_{i})$. Therefore, because of the non-negativity of mutual information, it is only possible that $I(z_{j};\bm{n}|y_{i})=0$ or, equivalently, $\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j}$.
Finally, we demonstrate that $(\bm{\tilde{y}}_{i}\leftrightarrow y_{i}\leftrightarrow z_{j})\,\land\,(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longrightarrow(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})$: We know by the Bayes’ theorem that $p(z_{j},\bm{y},\bm{n})=p(z_{j}|\bm{y},\bm{n})p(\bm{n}|\bm{y})p(\bm{y})=p(\bm{n}|z_{j},\bm{y})p(z_{j}|\bm{y})p(\bm{y})=p(\bm{n}|z_{j},\bm{y})p(z_{j}|y_{i})p(\bm{y})$. Thus, $p(z_{j}|x)=p(z_{j}|\bm{y},\bm{n})=p(z_{j}|y_{i})$ if and only if $p(\bm{n}|\bm{y})=p(\bm{n}|z_{j},\bm{y})$, which is equivalent to having $p(z_{j}|\bm{y})=p(z_{j}|\bm{y},\bm{n})=p(z_{j}|\bm{x})$ and since $p(z_{j}|\bm{y})=p(z_{j}|y_{i})$, we are left with $p(z_{j}|\bm{x})=p(z_{j}|y_{i})$, and thus, $(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})$. ∎
@@
Appendix B Training Hyperparameters
@@ Appendix C VAE Architectures
C.1 Architecture Following Burgess et al. (2018)
@@
C.2 Architecture Following Chen et al. (2018)
@@
C.3 Architecture Following Locatello et al. (2020)
@@
C.4 Architecture Combining Locatello et al. (2020) and Watters et al. (2019)
@@
C.5 Small Architecure Following Montero et al. (2022)
@@
C.6 Large Architecure Following Montero et al. (2022)

### v3

@@ Appendix A Proofs of Theorems
Let $\bm{y}=\{y_{l}\}_{l=1}^{n}$ denote a set of factors and $\bm{z}=\{z_{j}\}_{j=1}^{m}$ a representation. If $z_{j}$ is a minimal representation of $y_{i}$, it follows that $z_{j}$ is factors-invariant with respect to $y_{i}$ and nuisances-invariant. Equivalently, we have that:
$(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longrightarrow(\bm{\tilde{y}}_{i}\leftrightarrow y_{i}\leftrightarrow z_{j})\,\land\,(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})$
@@
Second, we demonstrate that $(\bm{x}\leftrightarrow y_{i}\leftrightarrow z_{j})\Longrightarrow(\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j})$: Given this Markov chain, we know from the DPI that $I(z_{j};\bm{x})\leq I(z_{j};y_{i})$. On the other hand, since $z_{j}$ is a representation of $\bm{x}$, we have the Markov chain $(\bm{n},y_{i})\leftrightarrow\bm{x}\leftrightarrow z_{j}$. Equivalently, we have that $I(z_{j};\bm{n},y_{i})\leq I(z_{j};\bm{x})$. The chain rule of mutual information tells us that $I(z_{j};\bm{n}|y_{i})=I(z_{j};\bm{n},y_{i})-I(z_{j}|y_{i})$. According to the above two points, we are left with $I(z_{j};\bm{n}|y_{i})\leq I(z_{j};\bm{x})-I(z_{j}|y_{i})$. Therefore, because of the non-negativity of mutual information, it is only possible that $I(z_{j};\bm{n}|y_{i})=0$ or, equivalently, $\bm{n}\leftrightarrow y_{i}\leftrightarrow z_{j}$. ∎
@@
Appendix B Influence of Number of Bins
B.1 Toy Experiment 1 (Section 5.1)
As observed in Figures 7 and 8, increasing the number of bins renders our metrics less sensitive to factor correlations ($\alpha$), while still consistently satisfying the expected behavior described in Section 5.1.
(a) 10 Bins
(b) 15 Bins
(c) 20 Bins
(d) 25 Bins
(e) 30 Bins
Figure 7: Comparison of Minimality for different number of bins (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest dependence, $\delta=1$).
(a) 10 Bins
(b) 15 Bins
(c) 20 Bins
(d) 25 Bins
(e) 30 Bins
Figure 8: Comparison of Sufficiency for different number of bins (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest dependence, $\delta=1$).
B.2 Toy Experiment 2 (Section 5.2)
In contrast to the previous section, Figure 9 reveals the opposite trend: increasing the number of bins makes Minimality more sensitive to $\beta$ (the amount of nuisances in the representation). Crucially, however, this increased sensitivity does not violate the expected behavior described in Section 5.2. On the other hand, Sufficiency remains invariant to the value of $\beta$, regardless of the number of bins used.
(a) Minimality
(b) Sufficiency
Figure 9: Comparison of Minimality and Sufficiency for different number of bins (y-axis) for different values of $\beta$ (x-axis)
B.3 Neural Representations Experiment (Section 6)
Figures 10 and 11 demonstrate that the rank correlations for key metric pairs are robust to discretization choices. Specifically, the correlation between Sufficiency and accuracy under low data and computational resources, and between Minimality and IRS, remain almost invariant to the number of bins used for calculation.
(a) DSprites
(b) MPI3D
(c) Shapes3D
Figure 10: Rank correlation between Sufficiency for different numbers of bins and accuracy of a random forest classifier trained on 1000 samples across different levels of factor dependence.
(a) DSprites
(b) MPI3D
(c) Shapes3D
Figure 11: Rank correlation between Minimality for different numbers of bins and Interventional Robustness Score across different levels of factor dependence.
Appendix C Influence of the Number of Samples
C.1 Toy Experiment 1 (Section 5.1)
As shown in Figures 12 and 13, the proposed metrics exhibit remarkable stability with respect to the number of samples, remaining consistent regardless of the factor correlations determined by $\alpha$.
(a) 1000 Samples
(b) 2000 Samples
(c) 5000 Samples
(d) 10000 Samples
Figure 12: Comparison of Minimality for different number of samples (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest dependence, $\delta=1$).
(a) 1000 Samples
(b) 2000 Samples
(c) 5000 Samples
(d) 10000 Samples
Figure 13: Comparison of Sufficiency for different number of samples (y-axis) across varying values of $\alpha$ (x-axis) and $\delta$ (color). Colors range from dark blue (highest dependence, $\delta=\frac{1}{n}$) to dark red (lowest dependence, $\delta=1$).
C.2 Toy Experiment 2 (Section 5.2)
As shown in Figure 14, the proposed metrics exhibit remarkable stability with respect to the number of samples, remaining consistent regardless of the level of nuisances determined by $\beta$.
(a) Minimality
(b) Sufficiency
Figure 14: Comparison of Minimality and Sufficiency for different number of samples (y-axis) for different values of $\beta$ (x-axis)
C.3 Neural Representations Experiment (Section 6)
Figures 15 and 16 demonstrate that the rank correlations for key metric pairs are robust to number of samples. Specifically, the correlation between Sufficiency and accuracy under low data and computational resources, and between Minimality and IRS, remain almost invariant to the number of samples.
(a) DSprites
(b) MPI3D
(c) Shapes3D
Figure 15: Rank correlation between Sufficiency for different numbers of samples and accuracy of a random forest classifier trained on 1000 samples across different levels of factor dependence.
(a) DSprites
(b) MPI3D
(c) Shapes3D
Figure 16: Rank correlation between Minimality for different numbers of samples and Interventional Robustness Score across different levels of factor dependence.
Appendix D Training Hyperparameters
@@ Appendix E VAE Architectures
E.1 Architecture Following Burgess et al. (2018)
@@
E.2 Architecture Following Chen et al. (2018)
@@
E.3 Architecture Following Locatello et al. (2020)
@@
E.4 Architecture Combining Locatello et al. (2020) and Watters et al. (2019)
@@
E.5 Small Architecure Following Montero et al. (2022)
@@
E.6 Large Architecure Following Montero et al. (2022)